A vertical spring has a force constant of 125N/m. By how much will the spring be compressed if a mass of 0.525kg is placed on top of it?


1
Expert's answer
2021-02-10T10:05:23-0500
"kx=mg,""x=\dfrac{mg}{k}=\dfrac{0.525\ kg\cdot9.8\ \dfrac{m}{s^2}}{125\ \dfrac{N}{m}}=0.041\ m."